The monthly income of employees in a small company is summarised in the frequency 

distribution below:

Monthly salary (N$) Number of employees

[1500 – 1999] 15

[2000 – 2499] 10

[2500 – 2999] 20

[3000 – 3499] 4

[3500 – 3999] 1

Required: 

a) Suppose the government regulations stipulate that the minimum monthly wage is N$ 2 500. What 

percentage of employees at this company earn below the stipulated minimum wage? (2)

b) How many employees earn exactly N$ 3 000 per month? (2)

c) How many employees earn at least N$ 3 000 per month? (1)

d) Calculate the following descriptive statistics for the random 


1
Expert's answer
2021-03-30T06:50:42-0400



(a) The number of people who earn below minimum age = 15+10=25


(b) The regression coefficient of y on x is-

"b_{yx}=\dfrac{n\sum xy-\sum x \sum y}{n\sum x^2-(\sum x)^2}"


"=\dfrac{5\times 129723.5-120475\times 50}{5\times 40298751.25-(120475)^2}"


"=\dfrac{-38757.5}{12500000}=-0.0031"


The regreesion equation of y on x is-

"y-\bar{y}=b_{yx}(x-\bar{x})"


"y-10=-0.0031(x-2749.5)"


Number of people Who earn exactly $3000


"y=10-0.0031(3000-2749.5)\\\Rightarrow y=10-0.0031(250)=10-0.77=9.22" ~=9


(c) Number of people who earn at least $3000

=4+1=5


(d) Mean of given data "\mu=\dfrac{\sum xf}{\sum f}=\dfrac{120475}{50}=2409.5"


Variance "\sigma^2 = \dfrac{\sum (x-\mu)^2}{N}=\dfrac{3078000}{50}=61560"


standard deviation "\sigma=\sqrt{variance}=\sqrt{61560}=248.11"